OpenStudy (anonymous):

What is the ratio of the velocity of sound in hydrogen to the velocity of sound in oxygen?!?

OpenStudy (anonymous):

hydrogen = 2/3 oxygen = 1/3 Assuming you're talking about the speed of sound in water

OpenStudy (mrnood):

What has this to do with water? The ratio of the 2 velocities is given by \[\frac{ SpeedHydrogen }{ SpeedOxygen }\] I certainly don't know those speed without looking them up, and I doubt whether anyone else her does. SO Look up the speeds in your text or on google Insert them into the above formula and calculate the result

OpenStudy (vincent-lyon.fr):

For an ideal gas, velocity is: \(v=\sqrt{\Large\frac{\gamma RT}{M}}\) Hydrogen and oxygen have the same \(\gamma\), because both are diatomic molecules, but have different molar masses.

OpenStudy (mrnood):

Using Vincent's formula I have come up with a number which is very close to the value derived from my formula and a list od speeds of sounds in gases. http://www.engineeringtoolbox.com/speed-sound-gases-d_1160.html